# When can isotropy help adapt LLMs’ next word prediction to numerical domains?

### v3

@@ 4 Study of isotropy in LLM hidden representations
Analysis settings. For illustrative purposes, in this section, we present two examples from a specific numerical domain, such as transport, and for a specific language model (i.e., GPT-2) that illustrate the conditions under which isotropy is preserved
Figure 2: Visualization of transport Dataset 1 (Left) and Dataset 2 (Right) labels for two extreme cases of velocities, i.e., $10$ km/hour and $100$ km/hour.
in LLM representations. In Section 5, we provide comprehensive evaluations across $22$ different datasets from various numerical domains and other $5$ different language models with different architectures and tokenization techniques. For this specific analysis, we select two datasets from transport as they are dynamic, noisy, time-varying, and thus hold all primary characteristics of the numerical data across various domains. Specifically, we use two ways of sharing signals which are typically used for signal transmission in transportation settings, as shown in Figure 2. We call these two datasets as “Dataset $1$” and “Dataset $2$”. The downstream task here is to predict the transport signalling property over wireless channel using LLM, where Dataset $1$ causes good downstream performance (i.e., near optimal signal prediction), while the Dataset $2$ causes bad downstream task performance (i.e., high error in signal prediction). We use NMSE as a performance metric for the numeric downstream task because it is widely used for signal prediction. We deploy the first six layers of GPT2 Radford et al. (2019) and use the datasets and simulation setups from Liu et al. (2024), which are the standard settings for time series forecasting. We predict $L=4$ future signaling properties based on the historical $T=16$ signal properties through time series forecasting using GPT2. The training and validation dataset contains $8,000$ and $1,000$ samples, respectively, with user velocities uniformly distributed between $10$ km/hour and $100$ km/hour. The test dataset contains ten velocities ranging from $10$ km/hour to $100$ km/hour, with $1,024$ samples for each velocity.
@@
Layer 1 2 3 4 5 6
Dataset 1 4 4 4 4 4 4
Dataset 2 1 1 1 1 1 1
@@
Effective Dimensions. In each layer of each model, we start with a data matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $|\mathcal{V}|$ represents the number of tokens in the input time series sequence, and $D$ corresponds to the embedding dimension. We apply PCA to reduce the dimensionality from $D$ to $m$ i.e., $\mathbf{\tilde{A}}\in\mathbb{R}^{|\mathcal{V}|\times m}$. Then the fraction of variance captured by the reduced representation is given by: $r_{m}=\frac{\sum_{i=0}^{m-1}\sigma_{i}}{\sum_{i=0}^{D-1}\sigma_{i}}$ where $\sigma_{i}$ denotes the $i$-th largest eigenvalue of the covariance matrix of $\mathbf{A}$. We define the $\epsilon$-effective dimension as $d(\epsilon)\triangleq\arg\min_{m}r_{m}\geq\epsilon$. For instance, if $d(0.8)=3$, it means that three principal dimensions retain $80$% of the variance. A higher $d$ suggests a more isotropic space Cai et al. (2021), where information is spread across multiple dimensions rather than being concentrated in a narrow subspace. Table 1 presents the values of $d(0.8)$ for different layers and models. Surprisingly, GPT-2 has so few effective dimensions, with $d(0.8)=4$ for Dataset $1$ and $d(0.8)=4$ for Dataset $1$, for layers 1 through 6, as compared to its original embedding dimensions $D=768$. The reduced dimensionality suggests that GPT-2’s embeddings lie in a subspace defined by a very narrow cone Ethayarajh (2019), and hence, there would be hardly any model capacity. Surprisingly, these language models are as successful as they are in numerical domains, given that most of their embedding vectors are as similar to one another as they are. These observations motivate us to look deeper into the contextual embedding space.
@@ 4.1 Clusters in the Contextual Embedding Space
Let $G(\mathbb{\Psi})\!=\!(g_{1}(\mathbb{\Psi}),\ldots,g_{|\mathcal{V}|}(\mathbb{\Psi}))^{\top}\!:\mathbb{R}^{|\mathcal{V}|\times D}\mapsto\mathbb{R}^{|\mathcal{V}|\times D}$ be the function for self-attention, i.e., $g_{i}(\mathbb{\Psi})=\text{softmax}(\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\Psi}^{\top})\mathbb{\Psi}$, where $\mathbb{\Lambda}\!=\!\mathbb{W}_{Q}\mathbb{W}_{K}^{\top}\in\mathbb{R}^{D\times D}$, and $\mathbb{W}_{Q}\!\!\in\!\!\mathbb{R}^{D\times m}$, $\mathbb{W}_{K}\!\!\in\!\!\mathbb{R}^{D\times m}$ are the parameter matrices for the query and key matrices of self-attention. The lemma below contributes to understanding why the isotropic property of pre-trained LLMs help to generalize to numerical domains. The proof of this lemma closely aligns with the approach in Kim et al. (2021). The proof of this lemma follows the analysis in Kim et al. (2021) is provided in Appendix B for completeness.
Lemma 4.1.
From Lemma 4.1, we can see that to minimize the norm of the gradient $\|\mathbb{J}\|_{2}$, we essentially need to make $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}$ small. When $\mathbb{\Lambda}$ is small and all the input time series token embeddings are centered at the origin, $\sum_{i=1}^{|\mathcal{V}|}\mathbb{\psi}_{i}\!=\!0$, we have $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}\right|^{2}\approx\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ (see Appendix B).
The theorem below shows that $\mathbb{\Lambda}$ minimizes the objective $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ and contains the $m$ largest eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of time series token embeddings, where $m$ is the rank of $\mathbb{\Lambda}$. The proof of Theorem 4.1 is provided in Appendix C.
Theorem 4.1.
Theorem 4.1 shows that the self-attention learns to perform a function for the numeric downstream tasks through training, which are closely related to the $m$ largest eigenvectors of LLM hidden representations. In other words, the self-attention mechanism effectively projects input time series tokens onto a low-dimensional contextual embedding space defined by the top eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$.
Figure 3: Variations in GPT-2’s hidden representation for different datasets from the same domain: (a) PCA plot of contextual embedding space for transport Dataset 1. (b) PCA plot of contextual embedding space for transport Dataset 2.
@@
Motivated by the findings from Lemma 4.1 and Theorem 4.1, we project the models’ hidden representations of into a lower-dimensional space by utilizing the $m\!=\!3$ largest eigenvectors through PCA, as shown in Figure 3. The three axes of the figure represent the first three principal components of the covariance matrix of LLM representations of each layer. For instance, in layer $6$, the first three principal components account for $76$% and $92$% of the total variance for Dataset $1$ and Dataset $2$, respectively. From Figure 3 a and Figure 3 b, we can see that there are disconnected or slightly overlapped islands that are far away from each other through layer 3 to layer 6. Note that the first principal dimension value spans from 00 to $2000$, significantly wider than the other $2$ dimensions, and dominates the total variance. A similar analogy can also be observed for transport Dataset 2 in Figure 3 b. In (1), the space isotropy is measured on pairs of arbitrary time series token representations, which could reside in two disconnected clusters. However, given that the variance is dominated by distances between clusters, such estimation would be biased by the inter-cluster distances. Hence, it is more meaningful to consider a per-cluster (i.e., local) investigation rather than a global estimate.
@@
We start by performing clustering on the LLM representations in the contextual embedding space. There are various methods for performing clustering, such as $k$-means, DBSCAN Ester et al. (1996). We select $K$-means clustering method because it is reasonably fast in high embedding dimensions. We use the classical silhouette score analysis Rousseeuw (1987) to determine the number of clusters $|C|$ in the contextual embedding space (see Appendix D for details). Since each LLM contextual embedding instance $\mathbb{\psi}(k_{i})$ belongs to a particular cluster through clustering, the cosine similarity should be measured after shifting the mean to the origin Mu and Viswanath (2018). Accordingly, we subtract the mean for each cluster (i.e., centroid) and calculate the adjusted $\zeta_{\text{cos}}$ in Section 2. Assuming we have a total of $|C|$ clusters, let $\mathbb{\psi}_{c}(k_{i})=\{\mathbb{\psi}_{c}^{1}(k_{i}),\mathbb{\psi}_{c}^{2}(k_{i}),\dots\}$ be the set of token $k$’s contextual embeddings in cluster $c\in C$, and $\mathbb{\psi}_{c}(k_{i})$ be one random sample in $\mathbb{\psi}_{c}(k_{i})$. We define the adjusted inter-token cosine similarity as
@@
Isotropy in Dataset 1. In this section, we provide an example of Dataset $1$ where a high prediction accuracy (i.e., low NMSE) is achieved by GPT2 based model. For instance, in Figure 4a, we compare the NMSE performance of our GPT2 based LLM with non-language models for different user velocities. From Figure 4 a, we can observe that the NMSE performance of all models gradually increased along with the increase in user velocity. This is because, with the increase in velocity, the transport signal characteristics rapidly change within a very short coherence time, resulting in increased prediction difficulty for the prediction model. The GPT2 based model consistently outperforms other baselines and demonstrates its high prediction accuracy.
Figure 4: (a) LLM outperforms all other baselines for all of the ten different velocities for Dataset 1. (b) Inter-type cosine similarities for Dataset $1$ with different velocities. $\zeta\textquoteright_{\textrm{cos}}$ are close to zero for all the layers, including layer 6, indicating that nearly perfect isotropy exists in the LLM embedding space for the Dataset $1$, which preserves the structure in the LLM’s hidden representations and causes the high prediction accuracy.
For illustrative purposes, we pick three user velocities: $10$ km/hour, $50$ km/hour, and $100$ km/hour, for isotropy assessment of Dataset $1$. The GPT2 based model achieves good NMSE performance for all of these three velocities, as shown in Figure 4a. We apply inter-type cosine similarity $\zeta^{\prime}_{\textrm{cos}}$ in (6) to measure the isotropy in GPT2 embedding space. From Figure 4b, we can see that the GPT2 has consistent near-zero cosine similarity values for all layers, including layer 6. This indicates that nearly perfect isotropy exists in the GPT2 embedding space for the Dataset $1$, which preserves the structure in the GPT2’s hidden representations and causes good downstream task performance.
Figure 5: (a) The NMSE performance of the LLM based model for Dataset 2 deteriorates significantly compared to Dataset 1. (b) Inter-type cosine similarities for Dataset $2$ for different velocities. Higher $\zeta_{\textrm{cos}}$ values indicate a weak isotropy (i.e., anisotropy) exists in the LLM embedding space which causes a lack of structure in the hidden representations, yielding poor prediction accuracy.
Isotropy in Dataset 2. In this section, we provide an example of Dataset $2$ where a poor prediction accuracy (i.e., high NMSE) is achieved by GPT2 based model. As shown in Figure 5 a, the NMSE performance fluctuates randomly for different velocities, while the NMSE performance for Dataset $1$ is gradually increasing with increase in the velocities. The NMSE performance for Dataset $2$ deteriorates significantly compared to the Dataset $1$. As before, with the three user velocities, the NMSE performance for Dataset $2$ for all of these velocities is worse as compared to Dataset $1$, as shown in Figure 5 a. From Figure 5 b, we can observe a weak isotropy (i.e., anisotropy) in the LLM embedding space for Dataset $2$, causing a lack of structure in the GPT2 hidden representations, and thus leading to bad downstream performance.
@@ 5 Experiments
Baselines. We consider popular pre-trained LLMs as the baselines for numeric downstream tasks, including PatchTST Nie et al. (2023), Lag-Llama Rasul et al. (2024), Moirai-1.0-R Woo et al. (2024), Chronos-T5 Ansari et al. (2024) and Choronos-Bolt (https://huggingface.co/autogluon/chronos-bolt-base). The models use different architectures, time series tokenization techniques and hyperparameters for numeric downstream tasks as summarized in Table 2.
@@
Datasets. We conduct a comprehensive evaluation using $12$ different real time series datasets from various numerical domains, including energy, nature, finance, healthcare, retail and transportation. The sources of these open-source datasets along with their descriptions, including how each dataset is used across different LLM can be found in Table 4 of Appendix E. We also illustrate our findings using KernelSynth Ansari et al. (2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ more synthetic datasets via Gaussian processes in Section 5. As in Section 4.1, we select two different datasets from each numerical domain (as shown in Table 3) and then perform qualitative analysis with synthetic datasets and quantitative analysis with real datasets. The results of both of this analysis are provided in the next two sections.

### v4

@@ 4 Study of isotropy in LLM hidden representations
Analysis Settings.
For this study, we consider five different language models including Chronos-T5 Ansari et al. (2024), Chronos-Bolt [^1] https://huggingface.co/autogluon/chronos-bolt-base, PatchTST Nie et al. (2023), Moirai-1.0-R Woo et al. (2024), and Lag-Llama Rasul et al. (2024). For illustration, we randomly select a real dataset (i.e., finance-Dataset 1) from a broader collection of 22 numerical datasets that we use in this paper since we see similar results with all of these datasets. The details of these models and datasets could be found in Section 5.
@@
Layer 1 2 3 4 5 6 7 8 9 10 11 12
Chronos-T5 4 4 4 4 4 4 4 4 4 4 4 4
Chronos-Bolt 1 1 1 1 1 1 1 1 1 1 1 1
PatchTST 1 1
Moirai 1 1 1 1 1 1
Lag-Llma 2 2 2 2 2 2 2 2
4.1 Effective Dimensions
@@
In each layer of each model, we start with a data matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $|\mathcal{V}|$ represents the number of tokens in the input time series sequence, and $D$ corresponds to the embedding dimension. We apply PCA to reduce the dimensionality from $D$ to $m$ i.e., $\mathbf{\tilde{A}}\in\mathbb{R}^{|\mathcal{V}|\times m}$. Then, the fraction of variance captured by the reduced representation is given by: $r_{m}=\frac{\sum_{i=0}^{m-1}\sigma_{i}}{\sum_{i=0}^{D-1}\sigma_{i}}$ where $\sigma_{i}$ denotes the $i$-th largest eigenvalue of the covariance matrix of $\mathbf{A}$. We define the $\epsilon$-effective dimension as $d(\epsilon)\triangleq\arg\min_{m}r_{m}\geq\epsilon$. For instance, if $d(0.8)=3$, then three principal dimensions retain $80$% of the variance. A higher $d$ suggests a more isotropic space Cai et al. (2021), where information is spread across multiple dimensions rather than being concentrated in a narrow subspace. Table 1 presents the values of $d(0.8)$ for different layers and models. Surprisingly, all of these models have very small effective dimensions as compared to original embedding dimensions. For instance, Chronos-Bolt has very small effective dimensions, with $d(0.8)=1$ for layers 1 through 12, as compared to its original embedding dimensions $D=512$. The small effective dimensionality is another way of telling that that Chronos-Bolts’s embedding vectors lie in a subspace defined by a very narrow cone Ethayarajh (2019), and consequently, their inter-token cosine similarity is large. If all the embedding vectors lie on a $1$-dimensional line, the inter-token cosine similarity would be close to $1$, and there would be hardly any model capacity. Surprisingly, despite having such low effective dimensionality, these language models still perform well in numerical domains. This counterintuitive result motivate us to look deeper into the contextual embedding space.
4.2 Spectral Alignment for Generalization in Numerical Settings
Let $G(\mathbb{\Psi})\!=\!(g_{1}(\mathbb{\Psi}),\ldots,g_{|\mathcal{V}|}(\mathbb{\Psi}))^{\top}\!:\mathbb{R}^{|\mathcal{V}|\times D}\mapsto\mathbb{R}^{|\mathcal{V}|\times D}$ be the function for self-attention, i.e., $g_{i}(\mathbb{\Psi})=\text{softmax}(\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\Psi}^{\top})\mathbb{\Psi}$, where $\mathbb{\Lambda}\!=\!\mathbb{W}_{Q}\mathbb{W}_{K}^{\top}\in\mathbb{R}^{D\times D}$, and $\mathbb{W}_{Q}\!\!\in\!\!\mathbb{R}^{D\times m}$, $\mathbb{W}_{K}\!\!\in\!\!\mathbb{R}^{D\times m}$ are the parameter matrices for the query and key matrices of self-attention. The lemma below provides insights into how the isotropic property of pre-trained LLMs enables generalization in numerical domains. The proof of this lemma follows the analysis in Kim et al. (2021) is provided in Appendix B for completeness.
Lemma 1.
From Lemma 1, we can see that, in onder to minimize the norm of the gradient $\|\mathbb{J}\|_{2}$, we essentially need to make $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}$ small. When $\mathbb{\Lambda}$ is small and all the input time series token embeddings are centered at the origin, $\sum_{i=1}^{|\mathcal{V}|}\mathbb{\psi}_{i}\!=\!0$, we have $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}\right|^{2}\approx\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ (see Appendix B).
Next, we prove that $\mathbb{\Lambda}$ minimizes the objective $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ and contains the $m$ largest eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of time series token embeddings, where $m$ is the rank of $\mathbb{\Lambda}$.
Theorem 2.
Proof.
The proof of Theorem 2 is provided in Appendix C. ∎
Theorem 2 shows that the self-attention mechanism effectively projects input time series tokens onto a low-dimensional contextual embedding space defined by the top eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$. This result reveals that the self-attention mechanism in LLMs implicitly aligns with the dominant directions (i.e., top eigenvectors) of the contextual embedding space, and hence, suggesting that isotropy is not just a geometric artifact but a learned structural property that supports effective generalization to numerical downstream tasks.
While the self-attention aligns input representations with the dominant eigenvectors of the embedding space, the alignment may vary across different subregions of the contextual embedding space due to variations in the input sequences, token types, or contextual patterns. As a result, the degree of isotropy may differ across subregions of the contextual embedding space, which motivates the need to assess isotropy at a local (i.e., cluster) level rather than relying solely on a global metric. The next section explores these local structural patterns and examines the geometry of the hidden representations through principal component analysis (PCA), which helps reveal how variance is distributed across embedding dimensions.
Figure 2: Isolated or slightly overlapping cluster islands exist in the contextual embedding space for all models. For brevity, we only show a few representative middle layers from each model.
@@ 4.3 Clusters in the Contextual Embedding Space
Motivated by the results of Lemma 1 and Theorem 2, this section investigates local structural patterns by projecting the LLMs’ hidden representations into a lower-dimensional space using the top $m\!=\!3$ eigenvectors via PCA, as shown in Figure 2. The three axes of the figure represent the first three principal components of the covariance matrix of LLM representations of each layer. For instance, in Figure 2b and 2d, the first three principal components account for $94$% of the total variance in layer 8 of Lag-Llama and $83$% in layer 2 of Moirai. From Figure 2 a, 2 b, 2 c and 2 d, we can see that there are disconnected or slightly overlapping islands that are far away from each other. In (1), the space isotropy is measured on pairs of arbitrary time series token representations, which could reside in two disconnected clusters. However, given that the variance is dominated by distances between clusters, such estimation would be biased by the inter-cluster distances. Hence, it is more reasonable to consider a per-cluster (i.e., local) investigation rather than a global estimate.
Isotropy within Clusters.
@@
We start by performing clustering on the LLM representations in the contextual embedding space. There are various methods for performing clustering, such as $K$-means and DBSCAN algorithm Ester et al. (1996). We select $K$-means clustering method because it is reasonably fast in high embedding dimensions. We use the classical silhouette score analysis Rousseeuw (1987) to determine the number of clusters $|C|$ in the contextual embedding space (see Appendix D for details). Since each LLM contextual embedding instance $\mathbb{\psi}(k_{i})$ belongs to a particular cluster through clustering, the cosine similarity should be measured after shifting the mean to the origin Mu and Viswanath (2018). Accordingly, we subtract the mean for each cluster (i.e., centroid) and calculate the adjusted $\zeta_{\text{cos}}$ in Section 2. Assuming we have a total of $|C|$ clusters, let $\mathbb{\psi}_{c}(k_{i})=\{\mathbb{\psi}_{c}^{1}(k_{i}),\mathbb{\psi}_{c}^{2}(k_{i}),\dots\}$ be the set of token $k$’s contextual embeddings in cluster $c\in C$, and $\mathbb{\psi}_{c}(k_{i})$ be one random sample in $\mathbb{\psi}_{c}(k_{i})$. We define the adjusted inter-token cosine similarity as
@@
To put it in a nutshell, this section provides a theoretical foundation showing that self-attention projects input tokens onto a low-dimensional subspace aligned with the dominant eigenvectors of the embedding correlation matrix. This alignment induces isotropy in LLM hidden representations, stabilizing the partition function and preserving the structure needed for reliable numerical downstream task performances. In Section 5, we extend this analysis by empirically evaluating how isotropy in different language models’ hidden representations correlates with time series forecasting performances across a wide range of numerical datasets, varying context lengths, and noise levels.
@@ 5 Experiments
Baselines. We consider popular pre-trained LLMs as the baselines for numerical downstream tasks, including Chronos-T5 Ansari et al. (2024) and Choronos-Bolt (https://huggingface.co/autogluon/chronos-bolt-base), PatchTST Nie et al. (2023), Moirai-1.0-R Woo et al. (2024) and Lag-Llama Rasul et al. (2024).
@@
The considered models use different architectures, time series tokenization techniques and hyperparameters for numerical downstream tasks. For instance, Lag-Llma use decoder only transformer, PatchTST and Moirai-1.0-R use vanilla Transformer enoder, while Chronos-T5 and Choronos-Bolt use encoder-decoder transformer. Different baselines achieve contextual embedding in different ways. For example, PatchTST focuses on tokenizing time series as patches and uses self-attention for modeling dependencies within each patch and across patches, while CHORNOS-T5 and CHRONOS-Bolt adapt language modeling architectures minimally and generate categorical tokens by applying scaling and quantization. The details of these baselines are summarized in Table 2.
@@
Datasets. We conduct a comprehensive evaluation using $12$ different real time series datasets from various numerical domains, including energy, nature, finance, healthcare, retail and transportation. The sources of these open-source datasets along with their descriptions, including how each dataset is used across different LLM can be found in Table 4 of Appendix E. We also illustrate our findings using KernelSynth Ansari et al. (2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ additional synthetic datasets via Gaussian processes in Section 5. We select two different datasets from each numerical domain (as shown in Table 3) and then perform qualitative analysis with synthetic datasets and quantitative analysis with real datasets. The results of these analyses are provided in the next two sections.
5.1 Qualitative Analysis
We now analyze the time series forecasting by the baseline LLMs qualitatively. We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and stochastic. We are particularly interested in the isotropic measurement in the LLM’s last layer as it is related to the logits and probabilistic inference as explained in Section 2. So all isotropic measure provided in this section is based on the last layer of the baselines.